OpenStudy (anonymous):

evaluate f(x) = -x^2 + 1 for x=-3 -4 -8 4 -9

OpenStudy (calculusxy):

\[\huge f(x) = -x^2+1\] \[\huge f(3) = (-3)^2 + 1\]

OpenStudy (calculusxy):

Evaluate the expression.

OpenStudy (drigobri):

The expression is f(x) = -x^2 + 1 For x=-3 So where has x, you replace it with -3 f(-3) = -(-3)² + 1 f(-3) = -(9) +1 f(-3) = -9 + 1 f(-3) = -8